OpenStudy (anonymous):

When 20 is divided by a positive integer n, the remainder is 2. How many possible values are there for n?

OpenStudy (campbell_st):

well n has to be less than 20 and not a factor of 20. so that rules out 1, 2, 4, 5, 10 and 20

OpenStudy (anonymous):

a little confused still?

OpenStudy (anonymous):

i am confused

OpenStudy (campbell_st):

well if n is larger than 20 the remainder will always be 20.... e.g 20/21 remainder 20 20/35 r = 20.... if n is a factor there is no remainder

OpenStudy (anonymous):

we can try them all 20 divided by 18 leaves a remainder of 2

OpenStudy (anonymous):

likewise 9

OpenStudy (anonymous):

also perhaps 6

OpenStudy (campbell_st):

likewise 3

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

think we got them all?

Directrix (directrix):

Let 20 = nk + 2 where k is an integer. 18= nk 18/n = k So n would have to be a factor of 18.

Directrix (directrix):

n could be 1, 2, 3, 6, 9, 18 (total of 6) but should we exclude 18?

OpenStudy (anonymous):

Oh, I see where you guys are getting at. Thanks! :)

OpenStudy (anonymous):

20/n = x + 2/n = (nx + 2) /n -> nx + 2 = 20 -> nx = 20 -2 = 18 The chance is: 1 * 18, 2 * 9, 3 * 6 Thus possible values of n = 1, 2, 3